November 22th 2024

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • CDet
      • Last TDC ribbon cable installed, going from bottom det freame too repeater
      • test new LV : all 3 supplies on hand and distribution box ready. All mmodule 36.6 A at 6.1 V

to get 5 V at NINO - will go with the current design

      • find GRINCH and hodoscope
      • started looking at DAQ crate : need to use the back connector for the 15 modules
      • triggers for CDet - pc
    • HCal
      • could take data while GEM not running - plan to move cables next week
      • d90
    • Beamline
      • 3 or more people helping with Cip for recabling after Thanksgiving - Paul will be on site late

2nd week of December

    • ECal
      • network switch online - patched fiber to router for VTP
      • will boot ECal crates and VTP this afternoon
      • breaker tripped : Mark moved a few crates on extension coord and balanced load on phase
      • cosmics firmware ready : will try with Simona with pulser when CODA is ready
    • GEMs
      • still moving platform
      • BB VTPs are up - need to boot 4th VTP
      • will get VTP setup in test lab to test histrogramming method
      • CPU in test lab is not booting. CPU might be bad, also possible issue with atedf3 - will troubleshoot this afternoon - hatest1
      • Bryaan will immplement MPD power trasnceiver check : willll bbMPD to him tst